OpenStudy (anonymous):

Find any points of discontinuity for the rational function y=x-8/x^2+6x-7

OpenStudy (tkhunny):

AS WRITTEN, x = 0, by trivial inspection. If you intended y = (x-8)/(x^2 + 6x - 7) = \(\dfrac{x-8}{x^{2} + 6x - 6}\), then that's a different story. Factor the denominator and see what drops out.